Panamint Systems Corporation is estimating activity costs associated with producing disk drives, tapes drives, and wire...

  1. Panamint Systems Corporation is estimating activity costs associated with producing disk drives, tapes drives, and wire drives. The indirect labor can be traced to four separate activity pools. The budgeted activity cost and activity base data by product are provided below.


    Activity Cost

    Activity Base
    Procurement $329,900 Number of purchase orders
    Scheduling 223,700 Number of production orders
    Materials handling 426,300 Number of moves
    Product development 762,500 Number of engineering changes
    Production 1,434,200 Machine hours

    Number of
    Purchase
    Orders

    Number of
    Production
    Orders

    Number
    of
    Moves

    Number of
    Engineering
    Changes


    Machine
    Hours

    Number
    of
    Units
    Disk drives 3,900    300       1,450 12         1,700 2,200
    Tape drives 1,500    165       540 8         10,000 3,700
    Wire drives 12,600    820       4,500 22         11,100 3,000

    The activity rate for the procurement activity cost pool is

    a.$62.90 per purchase order

    b.$18.33 per purchase order

    c.$256.73 per purchase order

    d.$174.09 per purchase order
Answer #1

Total purchases orders=(3900+1500+12600)=18000

Activity rate=Respective cost/Respective activity

=329,900/18000

which is equal to

=$18.33 per purchase order(Approx)